Hollow Spheres Part A Consider the following arrangement of two conducting hollow spheres with a point charge of Qo-4.90 HC at the center. The inner sphere has a radius of 0.011 m and carries a net charge of Q1- -1.30 uC. The outer sphere has a radius of 0.061m and carries a net charge of Q2 -6.30 HC Lo Calculate the magnitude of the electric field at point A located at a distance 0.021 m from the center. 7.34x107 N/C Part B Calculate the surface charge density on the outer surface of the outer sphere. Neglect the thickness of the sphere. 222 C/m*2 Submit Answer Incorrect. Tries 1/6 Previous Tries
Answer #1

given

Qo = 4.9 uC

Q1 = - 1.3 uC

Q2 = 6.3 uC

inner radius = 0.011 m

outer radius = 0.061 m = R

r = 0.021 m

a )

using equation

EA = k ( Qo + Q1 ) / r2

= 9 x 109 x ( 4.9 x 10-6 - 1.3 x 10-6 ) / 0.0212

EA = 7.3469 x 107 N/C

b )

\sigma = Q / A

= ( Qo + Q1 + Q2 ) / R2

= ( 4.9 - 1.3 + 6.3 ) x 10-6 / 0.0612

\sigma = 2.3945176 x 107 N/C